# Robust $\mathcal{H}_{\infty}$ control and worst-case search in constrained parametric space

### v5

@@ III-C2 Monte-Carlo (MC) sampling
$\mathrm{P}(\mathrm{P}(f(\delta)>f(\delta^{*})\leq\epsilon)\geq 1-\gamma$ (21)
where $\delta^{*}$ corresponds to the lowest value of the objective function $f$ (in our case: $a(\delta)$, $h_{\infty}(\delta)$ or $h_{2}(\delta)$) among the $N$ samples. Note that this bound is independent from the number of uncertain parameters, from their variation ranges or distributions. Then, a local search with SQP can be performed starting from this worst case.
A third option combines the two previous approaches. It consists of drawing $n$ groups of $N$ samples, and for each group, selecting the worst-case sample, denoted $\delta^{*}_{1},\delta^{*}_{2},...\delta^{*}_{n}$, to initialize a SQP search. Each of the $\delta^{*}_{i}$ verifies the probability of Eq. (21). Additionally, the minimum among all $\delta^{*}_{i}$ also verifies a condition of the same kind, but among $nN$ samples. This strategy enables the local search to be initialized from multiple points, thereby increasing the coverage of the parametric space compared to the second approach, while requiring fewer SQP runs (more costly than evaluating random points) than the first approach. This approach is applied in our numerical application in Section V-C.
IV Robust control in constrained parametric space
This Section addresses robust controller optimization i.e. problems (1) and (2) in parametric space $\mathcal{D}$ given in (5).
IV-A Robust controller optimization algorithm
The Algorithm 2, described in a form that solves the standard $\mathcal{H}_{\infty}$ control problem (1), follows the approach proposed in [3] that relies on iteratively constructing a subset $\mathcal{D}_{a}\subset\mathcal{D}$ of so-called active configurations as follows. The multi-objectives formulation is tackled in Section IV-B.
The set of active configurations is initialized with the nominal configuration. Step 1 optimizes the controller for all current active configurations, based on the nonsmooth optimization algorithm of [8] implemented in the routine systune. Step 2 addresses the worst-case search in stability as described in Section III. If a Karush-Kuhn-Tucker (KKT) point of the optimization problem is found that leads to a spectral abscissa greater than a tolerance $\alpha_{max}$, it is added to $\mathcal{D}_{a}$. It is also possible to identify multiple KKT points during Step 2. Similarly, Step 3 tackles the worst-case search in performance and adds to $\mathcal{D}_{a}$ the worst-case configuration(s) that degrade the performance by more than a tolerance $\epsilon$. Finally, Step 4 seeks to validate the design with a more exhaustive worst-case search, meaning that the optimization parameters are tuned to ensure a more thorough exploration of potential worst-case configurations. In contrast, in Steps 2 and 3, the parameters are chosen to prioritize fast convergence to KKT points; additional quantitative details are provided in the application of Section V-D. For Steps 2, 3 and 4, a global exploration is performed with Monte-Carlo (MC), and a subsequent Sequential Quadratic Programming (SQP) refines the search until finding a KKT point.
@@ IV-A Robust controller optimization algorithm
Step 2: Destabilization.
@@
If $\alpha(A(K^{*},\delta))>\alpha_{max}$, add $\delta^{*}$ to $\mathcal{D}_{a}$ and go back to step 1.
Otherwise continue to step 3.
Step 3: Degrade performance.
[…]
If a destabilizing $\delta$ is found during this optimization, go back to step 2 and start SQP directly from $\delta$.
[…]
Otherwise, go to step 4.
Step 4: Validation
Perform exhaustive worst-case search in stability and performance using MC+SQP.
If a configuration $\delta^{*}$ not respecting criteria of step 2 or 3 is found, add $\delta^{*}$ to $\mathcal{D}_{a}$ and go back to step 1.
Otherwise, terminate.
IV-B Multi-objective formulation
Algorithm 2 is presented to address the standard problem formulation (1) involving a single $\mathcal{H}_{\infty}$ criterion. We note that this formulation also directly addresses the multi-objective problem:
[…]
since it is equivalent to a single objective problem (e.g. as in [9, 10]):
[…]
In our numerical implementation, in this case, we used the expression on the left side of Eq. (22); indeed, it was faster to compute the norm of each requirement, than the norm of the block diagonal system.
More generally, we consider the multi-objectives problem (2). As explained in [4], the controller optimization [8] of Step 1 treats it as :
@@ IV-B Multi-objective formulation
where $\beta$ is the soft performance obtained in Step 1:
@@
i.e. the algorithm has converged when the worst-case search degrades the soft requirement by less than $1+\epsilon_{1}$, and the hard requirements are less than $1+\epsilon_{2}$. Finally, we point out that, in our implementation, the Monte-Carlo sampling computes all norms involved in Eq. (23) and determines their maximum; whereas the subsequent SQP only evaluates the transfer function that achieved this maximum.
V Illustration: a large flexible spacecraft
V-A Plant model and uncertain parameters
The benchmark used in this study is designed to be challenging regarding robustness. We consider a spacecraft composed of one rigid body connected to two large flexible solar panels of clamped-free natural frequencies between 0.20 and 0.85 rad/s. Uncertainties are considered on various mechanical properties; in total, there are 43 uncertain parameters $\delta_{i}$ (normalized such that $1\leq\delta_{i}\leq 1$, without loss of generality) regrouped in a vector $\delta$. Noting $M_{P}(\delta)$ the mass matrix at attachment point $P$ of a solar panel, and $L_{P}(\delta)$ the matrix of modal participation factors, the robustness analysis should exclude any configuration $\delta$ such that the residual mass matrix defined in Eq. (24) is not positive definite (indeed, such configurations do not represent a physical system). Hence the nonlinear constraint:
@@ V-A Plant model and uncertain parameters
$c(\delta):=\alpha(-M_{r}(\delta))\leq 0\;.$ (25)
The spacecraft model $G(s,\delta)$ is generated with the multibody approach detailed in [1] and implemented in the Satellite Dynamics Toolbox Library [2, 37]. The block $\Delta_{G}$ of the Linear Fractional Representation (LFR) model, such that $G(s,\delta)=\mathcal{F}_{u}(M_{G}(s),\Delta_{G})$ (upper LFT), has size 204. The LFR model of the spacecraft includes all configurations for $1\leq\delta_{i}\leq 1$ even though only $\sim 22\%$ of them respect the nonlinear constraint (25).
Remark 8.
The constraint (25) is not guaranteed to be differentiable everywhere, which contradicts the assumptions of Section 2. However, since this is the formulation that naturally arises in our application, we still test it numerically.
V-B Closed loop and control problem
The benchmark focuses on the attitude control of the satellite described above to follow a reference $r$. The control input $u$ of the plant $G(s,\delta)$ is the torque applied to the satellite, and the output vector contains the 3 angle measurements polluted by some noise $n$. The closed loop is represented in Fig. 2, where $d(s)$ represents actuator dynamics. We note $\mathcal{K}$ the set of controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, each one taking the measured angle and rate as inputs and generating a control torque, resulting in 24 decision parameters per controller (72 in total). The closed-loop plant has 50 states.
@@ V-B Closed loop and control problem
The control problem consists in placing a control bandwidth of 0.1 rad/s (hence interacting with the uncertain flexible modes and causing a strong sensitivity to the uncertain parameters) while ensuring a modulus margin of 0.5 and minimizing the variance of the actuator command in response to measurement noise. It reads:
@@
V-C Worst-case search
In this section, we present the worst-case search for the spectral abscissa and the $\mathcal{H}_{\infty}$ norm, as presented in Section III. We selected three controllers $K_{1}$, $K_{2}$, $K_{3}$ obtained at early iterations of the robust controller tuning presented in Section V-D and analyze the robustness of the closed-loop system with three Test cases. For the sake of brevity, and since all 3 Test cases yield similar conclusions, we present only the Test case 1 in the body of the paper; Test cases 2 and 3 are in Appendix B. Test case 1 illustrates how the proposed approaches perform in presence of several local optima, in different regions of the parametric space. We search for $\max_{\delta\in\mathcal{D}}\;\alpha(A(K_{1},\delta))$.
The tests are ran on a 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz, with 14 cores and 16GB of RAM. Computation times are provided for reference, but they may vary depending on hardware, software configuration, implementation details, optimization parameters, etc.
a) SQP
We ran the Sequential Quadratic Programming (SQP) algorithm from 100 random initial points. As termination criteria, we set $10^{-6}$ as optimality tolerance (KKT stationarity condition), and $10^{-8}$ as step tolerance (step size between two iterations). The resulting worst-case spectral abscissas $\alpha(A(\delta^{*}_{\text{SQP}}))=-a(\delta^{*}_{\text{SQP}})$ of each run are in Fig. 3(a), and suggest the presence of several local optima, which highlights the limitation of such a local method. These various local optima may correspond to significantly different parametric configurations. For example, the worst case 0.135 has the value -1.0 for the (normalized) value of one of the moments of inertia of the solar panel, -0.09 and 1.0 for the modal participation factors of the clamped-free modes 1 and 3 in rotation around the same axis; in contrast, the local optimum 0.007 has -0.30, -1.0 and -0.87 respectively for these three parameters. Excluding the runs that did not detect instability (each of which took less than 0.1s), the remaining runs took between 2.5 and 78s (average: 21s – regarding total computation time, we recall that each SQP run is inherently sequential, but that the 100 runs are executed in parallel).
(a) SQP, with 100 random initializations
(b) MC and MC+SQP, repeated 100 times
Figure 3: Worst-case spectral abscissa for Test case 1
b) MC, MC + SQP
As explained in Section III-C2, and to compare the results with SQP alone, we evaluated 100 groups of 11000 samples, and the worst case in each group was selected to initialize an SQP run. Fig. 3(b) presents the worst case for each group (in blue) and the worst case after the subsequent SQP (in red). Overall, MC+SQP performs signicantly better than SQP alone or MC alone. Each group of 11000 samples, computed in parallel over 14 cores, took about 6.2s; the SQP runtimes were between 40s and 100s.
We mention that 78% of the drawn samples were not admissible ($c(\delta)>0$) and were therefore discarded before evaluation of $a(\delta)$. We also mention that the total number of runs $N_{\text{MC}}$ verifies $N_{\text{MC}}\approx\textrm{ln}(\gamma)/\textrm{ln}(1-\epsilon)$ with the values $\gamma=\epsilon=$1e-5, which guarantees as in Eq. (21) that the worst-case $a(\delta^{*}_{\text{MC}})$ found through all samples of Fig. 3(b) (without considering the subsequent SQP) covers 99.999% of the probability distribution of $a(\delta)$ with a confidence level of 99.999%. These observations suggest that the worst case around 0.136 is rare, even though unstable configurations are frequent (around 92% of the admissible configurations).
As mentionned earlier, Appendix B presents two additional Test cases with similar conclusions.
V-D Robust control
In this section, we solve the robust control Problem (26) using Algorithm 2.
V-D1 Implementation details
[…]
Step 1: Multi-models controller synthesis
systune is initialized with the controller obtained at previous iteration along with 13 random initializations. If no satisfactory controller is found, the process is repeated with new random initializations until one is obtained. In principle, a maximum number of attempts could be enforced to declare the problem infeasible, but this option was not activated, as a feasible solution was always found in our experiments.
Steps 2 and 3: Worst-case searches
Compared to the choice of parameters of Section V-C, we reduced the step tolerance from 1e-8 to 1e-6 for the SQP algorithm. The objective is to speed up the detection of active configurations in most iterations, relying on Step 4 to perform a finer validation. Below, we present two different strategies for the worst-case searches.
Strategy 1 (single KKT point)
We run a single instance of MC+SQP, i.e. a single run of SQP chosen from the worst case among the $N$ MC samples. Although, as seen in Section V-C, this does not always converge to the global optimum, we prioritize computational speed by stopping at the first KKT point for several reasons (any missed problematic configuration may still be found in subsequent steps or iterations). The resulting (single) KKT point, if problematic, is added to the active configurations.
Strategy 2 (multiple KKT points)
In Strategy 2, we perform 4 instances of MC+SQP and obtain 4 KKT points $\delta^{*}_{i}$. For clarity, let us assume that they are all problematic configurations. The worst case among those 4, noted $\delta^{*}_{1}$, is added to the active configurations. Then, we compute the distance $|\delta_{2}^{*}-\delta_{1}^{*}|$; if the distance exceeds a specified threshold, $\delta^{*}_{2}$ is also added to the active configurations. The procedure is repeated for $\delta_{3}^{*}$ and $\delta_{4}^{*}$: each $\delta_{i}^{*}$ is added to the active scenarios only if its distance from all previously selected points exceeds the threshold. The threshold is chosen as half of the hypercube diagonal of the parameter space, i.e., $\sqrt{k}$, where $k$ is the number of uncertain parameters. This choice limits the number of active scenarios by enforcing a minimum spatial separation, thereby ensuring that only significantly distinct worst-case configurations are retained. Indeed, the computation time of the multi-models controller synthesis (Step 1) increases with the number of active configurations.
Step 4: Validation
In Step 4, we run a similar worst-case search as in Section V-C. We generated only 20 groups of MC samples (with one subsequent SQP run for each of the 20 groups), which is deemed sufficient for our application based on the results of Section V-C. This procedure is applied for both the stability and performance, and the Algorithm 2 terminates when no problematic configuration is found. In the implementation, we conducted the worst-case search for performance before that for stability to avoid unnecessary computations, since performance was observed to fail more frequently than stability.
V-D2 Results
As in section V-C, the tests are ran on a 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz, with 14 cores and 16GB of RAM, and computation times are provided for reference. We used the tolerances $\epsilon_{1}=5\%$ and $\epsilon_{2}=1\%$ (cf. Eq. (23)).
@@ V-D2 Results
The Algorithm 2 was executed 5 times for each strategy. All runs converged to a robust controller. Table I presents a few averaged metrics for comparison. The runtime of the final validation (Step 4) is averaged over all 10 runs, because this step is conducted identically for each strategy. Fig. 4 presents the transfer functions of Problem (26) after optimization of the controller. Active configurations (in red) include the worst cases detected in the final iteration.
TABLE I: Controller optimization results
Metrics Strategy 1 Strategy 2
Nb of iterations ( = nb. of executions of Step 1) (min/avg/max) 7/8.4/10 6/7.4/10
Nb. of active configurations (min/avg/max) 7/8.6/12
Total runtime, excluding final validation (avg) 4034s 3708s
… including % represented by controller tuning (Step 1) 34% 33%
@@
… including % represented by worst-case searches (Steps 2 and 3) 66% 67%
Runtime of the final validation (Step 4 – stability) (avg) 139s
Runtime of the final validation (Step 4 – performance) (avg) 2332s
@@
Figure 4: Transfer functions in Problem (26) after controller optimization, with active configurations and 100 random samples
VI Conclusion

### v6

@@ 3.3 Global exploration, local exploitation
$\mathrm{P}(\mathrm{P}(f(\delta)<f(\delta^{*})\leq\epsilon)\geq 1-\gamma$ (16)
where $\delta^{*}$ corresponds to the lowest value of the objective function $f$ (in our case: $a(\delta)$ or $h_{\infty}(\delta)$) among the $N$ samples, i.e. that $\delta^{*}$ covers a proportion $1-\epsilon$ of the distribution with a confidence of $1-\gamma$. Note that this bound is independent from the number of uncertain parameters, from their variation ranges or distributions. Then, a local search with SQP can be performed starting from this worst case to refine the result. Alternatively, it is also possible to draw $M$ independent groups of $N/M$ samples and to initialize a SQP with the worst case of each group; the stochastic guarantee is the same as in Eq. (16) (tied to the worst case among the total of $N$ samples) but it allows to perform multiple local searches from different points of the parametric space to increase coverage. If those SQP runs can be parallelized, this procedure comes at no extra cost.
4 Robust control in constrained parametric space
This Section leverages the worst-case search of previous section, to address robust controller optimization i.e. Problems (1) and (2) in the parametric space $\mathcal{D}$ given in (5).
4.1 Robust controller optimization algorithm
The Algorithm 2, described in a form that solves the standard $\mathcal{H}_{\infty}$ control problem (1), follows the approach proposed in [3]. It relies on iteratively constructing a subset $\mathcal{D}_{a}\subset\mathcal{D}$ of so-called active configurations and, in this paper, essentially exploits the novel worst-case search strategy introduced in previous sections. The multi-objectives formulation is tackled in Section 4.2 with minor modifications to Algorithm 2.
The set of active configurations is initialized with the nominal configuration. Step 1 optimizes the controller for all current active configurations, based on the nonsmooth optimization algorithm of [8] implemented in the routine systune. Step 2 addresses the worst-case search in stability as described in Section 3. If a Karush-Kuhn-Tucker (KKT) point $\delta^{*}$ of the optimization problem is found that leads to a spectral abscissa greater than a tolerance $\alpha_{max}$, it is added to $\mathcal{D}_{a}$. It is also possible to identify multiple KKT points during Step 2. Similarly, Step 3 tackles the worst-case search in performance and adds to $\mathcal{D}_{a}$ the worst-case configuration(s) that degrade the performance by more than a tolerance $\epsilon$. Finally, Step 4 seeks to validate the design with a more exhaustive worst-case search, meaning that the number of Monte-Carlo samples can typically be larger than in Steps 2 and  3 (where faster execution is privileged) to ensure better stochastic guarantees in accordance with Eq. (16).
@@ 4.1 Robust controller optimization algorithm
Step 2: Worst-case search in stability.
@@
If $\alpha(A(K^{*},\delta))>\alpha_{max}$, add $\delta^{*}$ to $\mathcal{D}_{a}$ and go back to step 1. Otherwise continue to step 3.
Step 3: Worst-case search in performance.
@@
Step 4: Validation
Perform a more exhaustive worst-case search in stability and performance using MC+SQP. If a configuration $\delta^{*}$ not respecting criteria of step 2 or 3 is found, add $\delta^{*}$ to $\mathcal{D}_{a}$ and go back to step 1. Otherwise, terminate.
4.2 Multi-objective formulation
Algorithm 2 is presented to address the standard problem formulation (1) involving a single $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ criterion. We note that this formulation also directly addresses the following multi-objective problem:
@@ 4.2 Multi-objective formulation
since it is equivalent to a single objective problem:
@@
More generally, we consider Problem (2). The controller optimization [8] of Step 1 treats it as described in in [4]. In our worst-case performance searches (in Steps 3 and 4), we use the normalization of Eq. (17), such that the performance degradation test of Step 3 becomes, in the multi-objectives case:
$\text{max \;}\left\{\frac{1}{(1+\epsilon_{1})\beta}\underset{\delta\in\mathcal{D}}{\text{max \;}}||T_{z_{1}w_{1}}(s,K,\delta)||\;,\right.\\ \left.\frac{1}{(1+\epsilon_{2})}\;\underset{2\leq i\leq n}{\text{max \;}}\underset{\delta\in\mathcal{D}}{\text{max \;}}||T_{z_{i}w_{i}}(s,K,\delta)||\right\}>1$ (17)
where $||\cdot||$ may refer to the $\mathcal{H}_{2}$ or $\mathcal{H}_{\infty}$ system norm, and where $\beta$ is the soft performance obtained in Step 1:
[…]
i.e. the algorithm has converged when the worst-case search degrades the soft requirement by less than $1+\epsilon_{1}$, and the hard requirements are less than $1+\epsilon_{2}$.
4.3 Adding one vs multiple KKT points to active set
Algorithm 2 gives the possibility to add one or multiple KKT points to the active configurations after a worst-case search is ran. This paragraph elaborates on how to select which KKT points to add when there are multiple available, e.g. after running several SQP in parallel.
Consider two KKT points $\delta_{1}^{*}$ and $\delta_{2}^{*}$. If their distance $\delta_{1}^{*}-\delta_{2}^{*}||_{2}$ is inferior to a predefined threshold, it is not necessary to add both of them. This threshold can typically be chosen as a fraction of the diagonal of the hypercube $2\sqrt{n}$, where $n$ is the number of uncertain parameters. However, this approach may not be very robust to the presence of uninfluential parameters, which may have very different values in $\delta_{1}^{*}$ and $\delta_{2}^{*}$ and thus contribute to the distance, without actually justifying to include both configurations to the active set.
We circumvent this limitation by weighting the distance with global sensitivity indices $w_{i}$, computed for each parameter. Such methods generally only require a few hundreds of samples, and thus come at limited extra cost. Then, the distance is defined as $W(\delta_{1}^{*}-\delta_{2}^{*})||_{2}$ with $W=\text{diag(}w_{i}$). The threshold may now be chosen as a fraction of 2w2||w|| the diagonal of the rescaled hyperrectangle. For example, in the numerical application of Section 5.4, we used the activity scores of [21], a global sensitivity metric derived from the active subspace method.
5 Illustration: control of an uncertain mechanical system
The benchmark used in this study is designed to be realistic in terms of complexity (number of states, uncertain parameters, and decision variables), challenging regarding robustness, and includes a nonlinear constraint on the uncertain parameters that is typical of mechanical systems. The tests are ran on a 12th Gen Intel(R) Core(TM) i7-12700H 2.30 GHz, with 14 cores and 16GB of RAM, on Matlab R2025a. Computation times are provided for reference, but they may vary depending on hardware, software configuration, implementation details, optimization parameters, etc.
5.1 Plant model and uncertain parameters
We consider a spacecraft, composed of one rigid body connected to two large flexible solar panels. Uncertainties are considered on various mechanical properties; in total, there are 43 uncertain parameters $\delta_{i}$. The spacecraft model $G(s,\delta)$ is generated with the multibody approach detailed in [1] and implemented in the Satellite Dynamics Toolbox of [2, 35]. The block $\Delta_{G}$ of the Linear Fractional Representation (LFR) model, such that $G(s,\delta)=\mathcal{F}_{u}(M_{G}(s),\Delta_{G})$ (upper LFT), has size 204. Noting $M_{P}(\delta)$ the mass matrix at attachment point $P$ of a solar panel, and $L_{P}(\delta)$ the matrix of modal participation factors, the robustness analysis should exclude any configuration $\delta$ such that the residual mass matrix defined in Eq. (18) is not positive definite (indeed, such configurations do not represent a physical system). Hence the nonlinear constraint:
@@ 5.1 Plant model and uncertain parameters
$c(\delta):=\alpha(-M_{r}(\delta))\leq 0$ (19)
and is assumed to be differentiable, at least close to worst-case configurations (which is often legitimate, as deficiency usually occurs along a single principal direction, leading to a simple maximum eigenvalue). The LFR model of the spacecraft includes all configurations for $1\leq\delta_{i}\leq 1$ even though only $\sim 22\%$ of them respect the nonlinear constraint (19).
5.2 Closed loop and control problem
The benchmark focuses on the attitude control of the satellite described above to follow a reference $r$. The control input $u$ of the plant $G(s,\delta)$ is the torque applied to the satellite, and the output vector contains the angle and rate measurements polluted by some noise $n$. The closed loop is represented in Fig. 2, where $d(s)$ represents actuator dynamics. We note $\mathcal{K}$ the set of controllers $K(s)$ of the desired structure: one 4-th order controller per spacecraft axis, resulting in 24 decision parameters per controller (72 in total). The closed-loop plant has 50 states.
@@ 5.2 Closed loop and control problem
The control problem consists in placing a control bandwidth of 0.1 rad/s while ensuring a modulus margin of 0.5 and minimizing the variance of the actuator command in response to measurement noise. It reads:
@@
5.3 Worst-case search
In this section, we illustrate the worst-case search for the spectral abscissa, as presented in Section 3. We selected a controller $K_{1}$ obtained at an initial iteration of the robust controller tuning presented in Section 5.4, and analyze the robustness of the closed-loop system. We search for $\max_{\delta\in\mathcal{D}}\;\alpha(A(K_{1},\delta))$.
SQP alone
We ran the SQP algorithm implemented in the Matlab routine fmincon [^1] As termination criteria, we set $10^{-6}$ as optimality tolerance (KKT stationarity condition), and $10^{-8}$ as step tolerance (step size between two iterations). from 100 random initial points. The resulting worst-case $\mathcal{H}_{\infty}$ norms of each run are in Fig. 3(a), and show the presence of several local optima, which highlights the limitation of such a local method. The runs took up to 51s (average: 4.7s).
Monte-Carlo + SQP
To show the benefit of initializing SQP after a Monte-Carlo search rather than from any random initial point, we repeated 100 times the method of Section 3 with $\epsilon=\gamma=0.01$: drawing $N=460$ samples, and selecting the worst-case point to initialize an SQP run. Fig. 3(b) presents the worst case for each group (in blue) and the worst case after the subsequent SQP (in red). Overall, MC+SQP performs consistently better than SQP alone or MC alone. Each group took about 0.57s (the samples were evaluated in parallel); the SQP runtimes were between 3s and 9s. Thus, the Monte-Carlo step comes at little to no extra cost (if parallel computing is available), and reduces the risk of falling into local minima.
(a) SQP
(b) MC and MC+SQP
Figure 3: Worst-case search results (spectral abscissa)
5.4 Robust control
In this section, we solve the robust control Problem (20) using Algorithm 2. We used the tolerances $\epsilon_{1}=5\%$ and $\epsilon_{2}=1\%$ (cf. Eq. (17)). The worst-case searches of Steps 2 and 3 were carried with $\epsilon=\gamma=0.01$ in Eq. (16) yielding $N=460$ Monte-Carlo samples, and Step 4 with $\epsilon=\gamma=0.001$ yielding $N=6910$ samples. We used $M=10$ (cf. the discussion following Eq. (16)).
@@ 5.4 Robust control
For testing purpose, since the algorithm involves randomness from both Monte Carlo sampling, Algorithm 2 was executed 10 times. All runs were initialized with the same nominal controller, and all converged to a robust controller. Table 1 presents a few metrics. In particular, the most costly phase is the controller optimization (Step 1). Fig. 4 presents the transfer functions of Problem (20) after optimization of the controller. Active configurations (in red) include the worst cases detected in the final iteration.
Table 1: Controller optimization results
Metrics Min. Average Max.
Nb of iterations 6 8.4 12
Nb of active configurations 12 16.3 24
Total runtime (s) 1312 2257 3954
… including Step 1* runtime 66% 74% 85%
@@
* Controller optimizaton (systune), as opposed to the worst-case searches (MC+SQP, Steps 2, 3 and 4)
@@
Figure 4: Transfer functions in Problem (20) after controller optimization, with active configurations (red), 100 random samples (blue) and requirement (dashed lines)
6 Conclusion